# In-medium bottomonium properties from lattice NRQCD calculations with extended meson operators

### v1

@@ I Introduction
Matsui and Satz proposed quarkonium suppression can serve as a signature of quark-gluon plasma (QGP) formation in heavy-ion collisions Matsui and Satz (1986). Their proposal was based on the observation that color screening within a deconfined medium can make the interaction between the heavy quark and anti-quark short ranged, leading to the dissolution of quarkonia in QGP. Since then the study of quarkonium production in heavy-ion collisions has been a large part of the experimental heavy-ion program; see, e.g., Refs. Aarts et al. (2017); Zhao et al. (2020) for recent reviews.
@@
The Matsui-Satz picture turned out to be incomplete. At nonzero temperature the potential is complex: in addition to the modification of its real part it also develops an imaginary part encoding the dissipative effects of the medium Laine et al. (2007); Brambilla et al. (2008). The modification of the real part of the potential also does not always correspond to an exponentially screened form Brambilla et al. (2008). In a weakly coupled quark-gluon plasma the imaginary part of the heavy quark potential leads to a thermal quarkonium yield, and in fact is the dominant source of the disappearance of different quarkonium states Escobedo and Soto (2008); Laine (2009). Recent lattice QCD studies of the complex potential at nonzero temperature do not find substantial evidence for color screening till distances $\lesssim 1$ fm, but find a large imaginary part Bala et al. (2022); Bazavov et al. (2024); Tang et al. (2024a). Irrespective of the detailed microscopic mechanism, we still expect all the quarkonium states to be dissociated (melted) at sufficiently high temperatures. This is consistent with the behavior of the spatial quarkonium correlation function at high temperatures Karsch et al. (2012); Bazavov et al. (2015); Petreczky et al. (2021).
@@
In-medium quarkonium properties are encoded in the corresponding spectral functions, which are related to the Euclidean time correlation functions calculable on the lattice. There have been many attempts to study in-medium properties of charmonium Umeda et al. (2005); Datta et al. (2003); Karsch et al. (2003); Datta et al. (2004); Asakawa and Hatsuda (2004); Jakovac et al. (2007); Ohno et al. (2011); Ding et al. (2012, 2018) and bottomonium Jakovac et al. (2007); Aarts et al. (2011a, b, 2013a, 2013b, 2014); Kim et al. (2015, 2018) in lattice QCD. These studies focused on in-medium modifications of ground states of S- and P-wave quarkonia. Previous lattice QCD studies of in-medium quarkonium have mostly used point meson operators, i.e., operators with quark and anti-quark fields located in the same spatial point, which are known to have non-optimal overlap with the quarkonium wave-functions, especially, with the excited states. As a result, these correlators are largely dominated by the vacuum continuum parts of the spectral function, and isolating the contributions from the in-medium bottomonium modification is very difficult Mocsy and Petreczky (2008); Petreczky et al. (2011); Burnier et al. (2015). Recently, the possibility of studying in-medium bottomonium properties using correlators of extended meson operators has been explored within the framework of non-relativistic QCD (NRQCD) Larsen et al. (2019, 2020a, 2020b). It was found that such operators have very good overlap with different S- and P-wave bottomonia, and therefore, the correlators are more sensitive to the in-medium bottomonium properties than the ones with point sources. Furthermore, these calculations found that the in-medium mass shifts are consistent with zero, but the thermal bottomonium widths are large Larsen et al. (2019, 2020a). This is consistent with the absence of screening in the real part of the potential up to $\lesssim 1$ fm distances, but a sizable imaginary part Shi et al. (2022); Tang et al. (2024b). In fact, the estimates of the width obtained in Refs. Larsen et al. (2019, 2020a) are larger than most phenomenological estimates; see Ref. Andronic et al. (2024). However, the above calculations have been performed on lattices with temporal extent $N_{\tau}=12$ and a single lattice spacing for any given temperature values. Therefore, it is very important to extend these calculations to larger temporal extent and having more than one lattice spacing per temperature value. Furthermore, the temperature region below $199$ MeV was not explored. In this paper, we present calculations of bottomonium correlation functions using NRQCD and extended meson operators in the temperature range $T=133-250$ MeV using lattices with temporal extent $N_{\tau}=16-30$ and two lattice spacings $a=0.0493$ and $0.0602$ fm.
@@
The remainder of the paper is organized as follows. In the next section, we discuss the lattice calculations of bottomonium correlators, including our lattice QCD setup. In Sec. III we discuss the determination of in-medium bottomonium masses and width. Finally, Sec. IV contains our conclusions. Some technical details, including the fitting processes to extract vacuum masses and in-medium parameters, are discussed in the appendix.
II The lattice NRQCD calculations of bottomonium correlators
@@
Because of the large bottom ($b$) quark mass, the bottomonium system can be described in terms of NRQCD, an effective field theory where the energy scale associated with the heavy quark mass is integrated out Caswell and Lepage (1986); Thacker and Lepage (1991). This effective field theory can be formulated on the lattice, and lattice NRQCD has been successfully used in studying bottomonium properties in vacuum Lepage et al. (1992); Davies et al. (1994); Meinel (2009); Hammant et al. (2011); Dowdall et al. (2012); Daldrop et al. (2012), since the large discretization effects associated with the large $b$ quark mass are avoided. NRQCD has also been used to study bottomonium properties at nonzero temperatures Aarts et al. (2011a, b, 2013a, 2013b, 2014); Kim et al. (2015, 2018); Larsen et al. (2019, 2020a). In this section, we describe our calculation of bottomonium correlators with extended meson operators at zero temperature and finite temperatures. Extended meson operators offer improved projection onto specific bottomonium states compared to point-like counterparts, making correlators less sensitive to the continuum Davies et al. (1994); Meinel (2009, 2010); Hammant et al. (2011); Dowdall et al. (2012); Daldrop et al. (2012); Larsen et al. (2019, 2020a). For this reason, the corresponding correlators are also more sensitive to the in-medium modification of bottomonium properties.
@@
In Sec. II.1 we discuss our lattice setup, including the extended bottomonium operators used in our study. Our results at zero temperature, including the bottomonium mass spectrum, are discussed in Sec. II.2. The temperature dependence of the bottomonium correlators is discussed in Sec. II.3.
II.1 The calculation details
@@
The lattice calculations in this work are carried out on background gauge fields including (2+1)-flavor dynamical sea quarks, which are implemented using the highly improved staggered quarks and the tree-level Symanzik gauge (HISQ/tree) action. The strange quark mass $m_{s}$ is fixed to its physical value, with light quark masses equal to $m_{s}/20$, corresponding to the Goldstone pion mass about 160 MeV, which is near the physical point. In this work, the spatial extents of the lattices are set to $N_{\sigma}=64$ and the temperature is varied by varying the temporal extent $N_{\tau}$ from $30$ to 16 at a fixed lattice spacing $a$. We perform calculations at values of the lattice gauge coupling $\beta=10/g_{0}^{2}=7.596$ and 7.373, corresponding to lattice spacings $a=0.0493$ and $a=0.0602$ fm, respectively. The physical value of the lattice spacing is fixed using the parametrization of the $r_{1}$ scale from Ref. Bazavov et al. (2018) and the value $r_{1}=0.3106(17)$ fm from Ref. Bazavov et al. (2010). The gauge configurations were generated using the software suite SIMULATeQCD Altenkort et al. (2022).
@@
With these lattice spacings and $N_{\tau}$ values, we cover the temperature range from 133 MeV to 250 MeV. We also perform calculations on $48^{3}\times 64$ and $64^{4}$ lattices for $a=0.0602$ and $a=0.0493$ fm, respectively, to provide a zero-temperature baseline, using gauge configurations generated by the HotQCD collaboration Bazavov et al. (2014).
@@
For the bottom quark on the lattice, the tree-level tadpole-improved NRQCD action including spin-dependent $v^{6}$ corrections is employed, i.e. exactly the same setup as in Refs. Meinel (2010); Larsen et al. (2019, 2020a). Here $v$ is the heavy-quark velocity inside quarkonium. The tadpole improvement is achieved by dividing the link variables by the tadpole parameter $u_{0}$. In our calculations, we take $u_{0}$ to be the fourth root of the averaged plaquette. The stability parameter $n$, which controls the time discretization steps in the evolution of heavy-quark propagator, is set to $n=16$. This value is larger than the ones used in Refs. Larsen et al. (2019, 2020a). Given the relatively small lattice spacings we found it is safer to use large values of $n$.
@@
The bottom quark mass, $M_{b}$, appearing in the NRQCD action, is fixed by demanding that the kinetic mass of $\eta_{b}$, defined as $E_{0}(M_{b},\mathbf{p})=\sqrt{\mathbf{p}^{2}+M_{\mathrm{kin}}^{2}(M_{b})}+\text{const}$, agrees with the value of the $\eta_{b}$ mass from the Particle Data Group. We use the values of $M_{b}$ determined in Larsen et al. (2019). For $\beta=7.596$, we performed an independent analysis of the kinetic mass and found a value of $aM_{b}=0.955(17)$, which is consistent with the previous result $aM_{b}=0.957(9)$ Larsen et al. (2019). The parameters of the lattice NRQCD calculations are summarized in Tab. 1.
II.2 Results in vacuum
@@
Figure 2: Comparisons of mass differences $\Delta M$ for different bottomonium states with respect to the 1S spin-averaged mass $\bar{M}_{1S}$ from lattice NRQCD calculations (points) with those from PDG Navas et al. (2024) (green solid lines). Green dashed line is the lattice-predicted value from Ref. Larsen et al. (2020a). Open points are obtained from analyses on 1S/1P (black), 2S/2P (blue) and 3S/3P (red) wave-function-optimized sources (circle for $a=0.0493$ fm and triangle for $0.0602$ fm), while filled ones from Gaussian sources.
II.3 Temperature dependence of bottomonium correlators
@@
III Bottomonium properties at nonzero temperature
@@
For the bottomonium system studied here, the spectral function is generally expected to contain several states below the open-bottom threshold, with numerous higher states forming a continuum at large frequencies. Thus, $\rho(\omega,T)$ can be decomposed into two parts: $\rho(\omega,T)=\rho_{\mathrm{med}}(\omega,T)+\rho_{\mathrm{cont}}(\omega)$. The continuum part, $\rho_{\mathrm{cont}}(\omega)$, is considered to be temperature independent, aligning with the above observations that the effective masses exhibit almost no temperature dependence at $\tau\lesssim$0.16 fm. On the other hand, the extended sources used in our work are designed to selectively enhance the overlap with particular states while suppressing contributions from undesired ones Larsen et al. (2020a, 2019); Wetzorke et al. (2002); Stickan et al. (2004). As a result, $\rho_{\mathrm{med}}(\omega,T)$ in vacuum can be parametrized as $\rho_{\mathrm{med}}(\omega,T=0)=A\delta(\omega-M)$, where $M$ represents the mass of a particular state targeted for projection. Combining these considerations, the continuum part can be isolated at zero temperature by $\rho_{\mathrm{cont}}(\omega)=\rho(\omega,T=0)-A\delta(\omega-M)$, leading to
@@
Here Eq. 1 is employed and $C_{\mathrm{cont}}(\tau)$ is defined as the Laplace transform of $\rho_{\mathrm{cont}}(\omega)$. The parameters $A$ and $M$ are extracted by fitting relevant correlators in vacuum, implemented as outlined in Appendix A. Subsequently, a continuum-subtracted correlator can be defined as in Refs. Larsen et al. (2020a, 2019),
@@
III.1 Continuum-subtracted effective masses and in-medium bottomonium properties from simple ansätze
@@
At higher temperatures we see that the continuum-subtracted effective masses decrease linearly in $\tau$ if $\tau$ is not too large. For $\tau$ values close to $1/T$ the continuum-subtracted effective masses show a sudden drop in many cases. It was suggested that the linear decrease of the continuum-subtracted effective masses is related to the width of the dominant peak in $\rho_{\rm med}(\omega,T)$ Larsen et al. (2019, 2020a). If $\rho_{\rm med}(\omega,T)$ would consist of a single Gaussian peak,
@@
then the slope of the continuum-subtracted effective masses would be given by $\Gamma_{G}^{2}$. Note, however, that a Gaussian form of the spectral function is not the only one that could result in nearly linear decrease of the continuum-subtracted effective masses, as discussed below. In NRQCD, there is an additional contribution to the spectral function well below the dominant peak, which corresponds to the forward-propagating heavy $q\bar{q}$ pair interacting with the background of backward-propagating light degrees of freedom from the heat bath Bala et al. (2022). This contribution will depend on the choice of the meson operator, as different meson operators may have different overlap with the multi-particle states Bala et al. (2022). From Fig. 5 we see that the rapid drop-off in the continuum-subtracted effective masses is different for Gaussian and wave-function-optimized sources, as expected. The slope of the linear decrease appears to be the same for both type of sources. Furthermore, the slope of the continuum-subtracted effective masses is larger for the $\chi_{b0}$(1P) state than for the $\Upsilon$(1S) state, possibly suggesting that the in-medium width for the $\chi_{b0}$(1P) state is larger than the one for the $\Upsilon$(1S) state. The differences between the continuum-subtracted effective masses computed using Gaussian and wave-function-optimized sources arise also due to the additional contributions from the regions of the respective spectral functions below the dominant peaks. As discussed below, these differences in the continuum-subtracted effective masses do not lead to differences in the in-medium bottomonium masses and widths. In addition, the behavior of the continuum-subtracted effective masses is influenced by the detailed choice of the wave-function-optimized sources, e.g. by the choice of $\tau_{0}$ in the generalized eigenvalue problem, without affecting the in-medium bottomonium parameters.
@@
To summarize, the continuum-subtracted effective masses indicate that different bottomonium states may exist above the chiral crossover temperature but may acquire a thermal width. The continuum-subtracted effective masses show a linear decrease at moderate $\tau$ values that can be possibly related to the width of the bottomonium states. The slope of this linear decrease of the continuum-subtracted effective masses with $\tau$ becomes larger with increasing temperature and is also larger for higher-lying bottomonium states. This may indicate that higher-lying bottomonium states, which are larger in size, are more affected by the medium, and thus have larger thermal width, and the thermal width increases with the temperature. The slope of the linear increase of the continuum-subtracted effective masses is independent of the choice of the meson operators (Gaussian or wave-function-optimized), and therefore likely corresponds to a physical effect, e.g., thermal broadening of the peak. We also observe a sudden drop in the continuum-subtracted effective masses around $\tau\sim 1/T$, which is not related to the in-medium modification of bottomonium but to the contribution to the spectral function well below the dominant peak.
To obtain information on the in-medium properties of different bottomonium states, we need an ansätze for $\rho_{\rm med}(\omega,T)$, and, in particular, a parametrization of the dominant peak. It is expected that spectral functions at nonzero temperature have a quasi-particle peak described by a Lorentzian form in the vicinity of the peak. This is confirmed by explicit calculations of the bottomonium spectral functions in the T-matrix approach Tang et al. (2024b). Furthermore, the spectral functions corresponding to wave-function-optimized operators in the T-matrix approach can be described by a Lorentzian form around the peak, and decay exponentially far away from the peak Tang et al. (2024b). As an example, in Fig. 8 we show the spectral function corresponding to the wave-function-optimized correlation function of the $\chi_{b}$(1P) state at $T=182$ MeV from the T-matrix calculations. The spectral function of a static quark-antiquark pair at high temperature calculated using hard thermal loop perturbation theory also has a similar form Burnier and Rothkopf (2013). Thus, the quasi-particle peak in the spectral function can be described by the following form:
@@
with $\Theta$ being the Heaviside step function. We call this form of the spectral function the cut Lorentzian form. Such a form of the spectral function has been used in the analysis of the static quark-antiquark correlators at nonzero temperatures Bazavov et al. (2024). It was shown that this form of the spectral function also leads to a nearly linear behavior of the continuum-subtracted effective masses in $\tau$ Bazavov et al. (2024). In practice, a possible parametrization of $\Theta_{\delta}$ could be the following:

### v2

# D-33615 Bielefeld, Germany

@@ 1 Introduction
Matsui and Satz proposed quarkonium suppression can serve as a signature of quark-gluon plasma (QGP) formation in heavy-ion collisions Matsui:1986dk . Their proposal was based on the observation that color screening within a deconfined medium can make the interaction between the heavy quark and anti-quark short ranged, leading to the dissolution of quarkonia in QGP. Since then the study of quarkonium production in heavy-ion collisions has been a large part of the experimental heavy-ion program; see, e.g., Refs. Aarts:2016hap ; Zhao:2020jqu for recent reviews.
@@
The Matsui-Satz picture turned out to be incomplete. At nonzero temperature the potential is complex: in addition to the modification of its real part it also develops an imaginary part encoding the dissipative effects of the medium Laine:2006ns ; Brambilla:2008cx . The modification of the real part of the potential also does not always correspond to an exponentially screened form Brambilla:2008cx ; specifically, for interquark distances less than the Debye screening length, the potential is altered by power-law thermal effects rather than exponential screening. Indeed, while color screening is well-established at high temperature in terms of the free energy of a static quark anti-quark pair through lattice calculations Bazavov:2018wmo , its manifestation in the real part of the heavy quark potential is still inconclusive. In a weakly coupled quark-gluon plasma the imaginary part of the heavy quark potential leads to a thermal quarkonium yield, and in fact is the dominant source of the disappearance of different quarkonium states Escobedo:2008sy ; Laine:2008cf . Recent lattice QCD studies of the complex potential at nonzero temperature, assuming a dominant spectral peak, do not find substantial evidence for color screening till distances $\lesssim 1$ fm, contrary to screened real part reported by studies based on Hard-Thermal-Loop inspired ansatz, although all find a large imaginary part Bala:2021fkm ; Bazavov:2023dci ; Tang:2023tkm . Irrespective of the detailed microscopic mechanism, we still expect all the quarkonium states to be dissociated (melted) at sufficiently high temperatures. This is consistent with the behavior of the spatial quarkonium correlation function at high temperatures Karsch:2012na ; Bazavov:2014cta ; Petreczky:2021zmz .
@@
In-medium quarkonium properties are encoded in the corresponding spectral functions, which are related to the Euclidean time correlation functions calculable on the lattice. There have been many attempts to study in-medium properties of charmonium Umeda:2002vr ; Datta:2002ck ; Karsch:2002wv ; Datta:2003ww ; Asakawa:2003re ; Jakovac:2006sf ; Ohno:2011zc ; Ding:2012sp ; Ding:2017std and bottomonium Jakovac:2006sf ; Aarts:2010ek ; Aarts:2011sm ; Aarts:2012ka ; Aarts:2013kaa ; Aarts:2014cda ; Kim:2014iga ; Kim:2018yhk in lattice QCD. These studies focused on in-medium modifications of ground states of S- and P-wave quarkonia. Previous lattice QCD studies of in-medium quarkonium have mostly used point meson operators, i.e., operators with quark and anti-quark fields located in the same spatial point, which are known to have non-optimal overlap with the quarkonium wave-functions, especially, with the excited states. As a result, these correlators are largely dominated by the vacuum continuum parts of the spectral function, and isolating the contributions from the in-medium bottomonium modification is very difficult Mocsy:2007yj ; Petreczky:2010tk ; Burnier:2015tda . Recently, the possibility of studying in-medium bottomonium properties using correlators of extended meson operators has been explored within the framework of non-relativistic QCD (NRQCD) Larsen:2019bwy ; Larsen:2019zqv ; Larsen:2020rjk . It was found that at zero temperature such operators have very good overlap with different S- and P-wave bottomonia Davies:1994mp ; Meinel:2009rd ; Meinel:2010pv ; Hammant:2011bt ; HPQCD:2011qwj ; Daldrop:2011aa while remaining unaffected by the continuum, and therefore, their correlators are expected to be more sensitive to in-medium bottomonium properties than the ones with point sources. Furthermore, at finite temperatures these extended operators are anticipated to preserve their utility, as the short-distance physics is largely unaltered by the thermal effects. Accordingly, these calculations at finite temperatures found that the in-medium mass shifts are consistent with zero, but the thermal bottomonium widths are large Larsen:2019bwy ; Larsen:2019zqv . This is consistent with the absence of screening in the real part of the potential up to $\lesssim 1$ fm distances, but a sizable imaginary part Shi:2021qri ; Tang:2024dkz . In fact, the estimates of the width obtained in Refs. Larsen:2019bwy ; Larsen:2019zqv are larger than most phenomenological estimates; see Ref. Andronic:2024oxz . However, the above calculations have been performed on lattices with temporal extent $N_{\tau}=12$ and a single lattice spacing for any given temperature values. Therefore, it is very important to extend these calculations to larger temporal extent and having more than one lattice spacing per temperature value. Furthermore, the temperature region below $199$ MeV was not explored. In this paper, we present calculations of bottomonium correlation functions using NRQCD and extended meson operators in the temperature range $T=133-250$ MeV using lattices with temporal extent $N_{\tau}=16-30$ and two lattice spacings $a=0.0493$ and $0.0602$ fm.
@@
The remainder of the paper is organized as follows. In the next section, we discuss the lattice calculations of bottomonium correlators, including our lattice QCD setup. In Sec. 3 we discuss the determination of in-medium bottomonium masses and width. Finally, Sec. 4 contains our conclusions. Some technical details, including the fitting processes to extract vacuum masses and in-medium parameters, are discussed in the appendix.
2 The lattice NRQCD calculations of bottomonium correlators
@@
Because of the large bottom ($b$) quark mass, the bottomonium system can be described in terms of NRQCD, an effective field theory where the energy scale associated with the heavy quark mass is integrated out Caswell:1985ui ; Thacker:1990bm . This effective field theory can be formulated on the lattice, and lattice NRQCD has been successfully used in studying bottomonium properties in vacuum Lepage:1992tx ; Davies:1994mp ; Meinel:2009rd ; Hammant:2011bt ; HPQCD:2011qwj ; Daldrop:2011aa , since the large discretization effects associated with the large $b$ quark mass encountered in the relativistic formulation are avoided. NRQCD has also been used to study bottomonium properties at nonzero temperatures Aarts:2010ek ; Aarts:2011sm ; Aarts:2012ka ; Aarts:2013kaa ; Aarts:2014cda ; Kim:2014iga ; Kim:2018yhk ; Larsen:2019bwy ; Larsen:2019zqv . In this section, we describe our calculation of bottomonium correlators with extended meson operators at zero temperature and finite temperatures. Extended meson operators offer improved projection onto specific bottomonium states compared to point-like counterparts, making correlators less sensitive to the continuum Davies:1994mp ; Meinel:2009rd ; Meinel:2010pv ; Hammant:2011bt ; HPQCD:2011qwj ; Daldrop:2011aa ; Larsen:2019bwy ; Larsen:2019zqv . For this reason, the corresponding correlators are also more sensitive to the in-medium modification of bottomonium properties.
@@
In Sec. 2.1 we discuss our lattice setup, including the extended bottomonium operators used in our study. Our results at zero temperature, including the bottomonium mass spectrum, are discussed in Sec. 2.2. The temperature dependence of the bottomonium correlators is discussed in Sec. 2.3.
2.1 The calculation details
@@
The lattice calculations in this work are carried out on background gauge fields including (2+1)-flavor dynamical sea quarks, which are implemented using the highly improved staggered quarks and the tree-level Symanzik gauge (HISQ/tree) action. The strange quark mass $m_{s}$ is fixed to its physical value, with light quark masses equal to $m_{s}/20$, corresponding to the Goldstone pion mass about 160 MeV, which is near the physical point. In this work, the spatial extents of the lattices are set to $N_{\sigma}=64$ and the temperature is varied by varying the temporal extent $N_{\tau}$ from $30$ to 16 at a fixed lattice spacing $a$. We perform calculations at values of the lattice gauge coupling $\beta=10/g_{0}^{2}=7.596$ and 7.373, corresponding to lattice spacings $a=0.0493$ and $a=0.0602$ fm, respectively. The physical value of the lattice spacing is fixed using the parametrization of the $r_{1}$ scale from Ref. Bazavov:2017dsy and the value $r_{1}=0.3106(17)$ fm from Ref. MILC:2010hzw . The gauge configurations were generated using the software suite SIMULATeQCD Altenkort:2021cvg .
@@
With these lattice spacings and $N_{\tau}$ values, we cover the temperature range from 133 MeV to 250 MeV. We also perform calculations on $48^{3}\times 64$ and $64^{4}$ lattices for $a=0.0602$ and $a=0.0493$ fm, respectively, to provide a zero-temperature baseline, using gauge configurations generated by the HotQCD collaboration HotQCD:2014kol .
@@
For the bottom quark on the lattice, the tree-level tadpole-improved NRQCD action including spin-dependent $v^{6}$ corrections is employed, i.e. exactly the same setup as in Refs. Meinel:2010pv ; Larsen:2019bwy ; Larsen:2019zqv . Here $v$ is the heavy-quark velocity inside quarkonium. The tadpole improvement is achieved by dividing the link variables by the tadpole parameter $u_{0}$. In our calculations, we take $u_{0}$ to be the fourth root of the averaged plaquette. The stability or Lepage parameter $n$, which controls the time discretization steps in the evolution of heavy-quark propagator, is set to $n=16$. This value is larger than the ones used in Refs. Larsen:2019bwy ; Larsen:2019zqv . Given the relatively small lattice spacings we found it is safer to use large values of $n$.
@@
The bottom quark mass, $M_{b}$, appearing in the NRQCD action, is fixed by demanding that the kinetic mass of $\eta_{b}$, defined as $E_{0}(M_{b},\mathbf{p})=\sqrt{\mathbf{p}^{2}+M_{\mathrm{kin}}^{2}(M_{b})}+\text{const}$, agrees with the value of the $\eta_{b}$ mass from the Particle Data Group. We use the values of $M_{b}$ determined in Larsen:2019bwy . For $\beta=7.596$, we performed an independent analysis of the kinetic mass and found a value of $aM_{b}=0.955(17)$, which is consistent with the previous result $aM_{b}=0.957(9)$ Larsen:2019bwy . The parameters of the lattice NRQCD calculations are summarized in Tab. 1.
2.2 Results in vacuum
@@
Figure 2: Comparisons of mass differences $\Delta M$ for different bottomonium states with respect to the 1S spin-averaged mass $\bar{M}_{1S}$ from lattice NRQCD calculations (points) with those from PDG ParticleDataGroup:2024cfk (green solid lines). Green dashed line is the lattice-predicted value from Ref. Larsen:2019zqv . Open points are obtained from analyses on 1S/1P (black), 2S/2P (blue) and 3S/3P (red) wave-function-optimized sources (circle for $a=0.0493$ fm and triangle for $0.0602$ fm), while filled ones from Gaussian sources.
2.3 Temperature dependence of bottomonium correlators
@@
3 Bottomonium properties at nonzero temperature
@@
For the bottomonium system studied here, the spectral function is generally expected to contain several states below the open-bottom threshold, with numerous higher states forming a continuum at large frequencies. Thus, $\rho(\omega,T)$ can be decomposed into two parts: $\rho(\omega,T)=\rho_{\mathrm{med}}(\omega,T)+\rho_{\mathrm{cont}}(\omega)$. The continuum part, $\rho_{\mathrm{cont}}(\omega)$, is considered to be temperature independent, aligning with the above observations that the effective masses exhibit almost no temperature dependence at $\tau\lesssim$0.16 fm. On the other hand, the extended sources used in our work are designed to selectively enhance the overlap with particular states while suppressing contributions from undesired ones Larsen:2019zqv ; Larsen:2019bwy ; Wetzorke:2001dk ; Stickan:2003gh . As a result, $\rho_{\mathrm{med}}(\omega,T)$ in vacuum can be parametrized as $\rho_{\mathrm{med}}(\omega,T=0)=A\delta(\omega-M)$, where $M$ represents the mass of a particular state targeted for projection. Combining these considerations, the continuum part can be isolated at zero temperature by $\rho_{\mathrm{cont}}(\omega)=\rho(\omega,T=0)-A\delta(\omega-M)$, leading to
@@
Here Eq. 1 is employed and $C_{\mathrm{cont}}(\tau)$ is defined as the Laplace transform of $\rho_{\mathrm{cont}}(\omega)$. The parameters $A$ and $M$ are extracted by fitting relevant correlators in vacuum, implemented as outlined in Appendix A. Subsequently, a continuum-subtracted correlator can be defined as in Refs. Larsen:2019zqv ; Larsen:2019bwy ,
@@
3.1 Continuum-subtracted effective masses and in-medium bottomonium properties from simple ansätze
@@
At higher temperatures we see that the continuum-subtracted effective masses decrease linearly in $\tau$ if $\tau$ is not too large. For $\tau$ values close to $1/T$ the continuum-subtracted effective masses show a sudden drop in many cases. It was suggested that the linear decrease of the continuum-subtracted effective masses is related to the width of the dominant peak in $\rho_{\rm med}(\omega,T)$ Larsen:2019bwy ; Larsen:2019zqv . If $\rho_{\rm med}(\omega,T)$ would consist of a single Gaussian peak,
@@
then the slope of the continuum-subtracted effective masses would be given by $\Gamma_{G}^{2}$. Note, however, that a Gaussian form of the spectral function is not the only one that could result in nearly linear decrease of the continuum-subtracted effective masses, as discussed below. In NRQCD, there is an additional contribution to the spectral function well below the dominant peak, which corresponds to the forward-propagating heavy $q\bar{q}$ pair interacting with the background of backward-propagating light degrees of freedom from the heat bath Bala:2021fkm . This contribution will depend on the choice of the meson operator, as different meson operators may have different overlap with the multi-particle states Bala:2021fkm . From Fig. 5 we see that the rapid drop-off in the continuum-subtracted effective masses is different for Gaussian and wave-function-optimized sources, as expected. The slope of the linear decrease appears to be the same for both type of sources. Furthermore, the slope of the continuum-subtracted effective masses is larger for the $\chi_{b0}$(1P) state than for the $\Upsilon$(1S) state, possibly suggesting that the in-medium width for the $\chi_{b0}$(1P) state is larger than the one for the $\Upsilon$(1S) state. The differences between the continuum-subtracted effective masses computed using Gaussian and wave-function-optimized sources arise also due to the additional contributions from the regions of the respective spectral functions below the dominant peaks. As discussed below, these differences in the continuum-subtracted effective masses do not lead to differences in the in-medium bottomonium masses and widths. In addition, the behavior of the continuum-subtracted effective masses is influenced by the detailed choice of the wave-function-optimized sources, e.g. by the choice of $\tau_{0}$ in the generalized eigenvalue problem, without affecting the in-medium bottomonium parameters.
@@
To summarize, the continuum-subtracted effective masses indicate that different bottomonium states may exist above the chiral crossover temperature but may acquire a thermal width. The continuum-subtracted effective masses show a linear decrease at moderate $\tau$ values that can be possibly related to the width of the bottomonium states. The slope of this linear decrease of the continuum-subtracted effective masses with $\tau$ becomes larger with increasing temperature and is also larger for higher-lying bottomonium states. This may indicate that higher-lying bottomonium states, which are larger in size, are more affected by the medium, and thus have larger thermal width, and the thermal width increases with the temperature. The slope of the linear decrease of the continuum-subtracted effective masses is independent of the choice of the meson operators (Gaussian or wave-function-optimized), and therefore likely corresponds to a physical effect, e.g., thermal broadening of the peak. We also observe a sudden drop in the continuum-subtracted effective masses around $\tau\sim 1/T$, which is not related to the in-medium modification of bottomonium but to the contribution to the spectral function well below the dominant peak.
To obtain information on the in-medium properties of different bottomonium states, we need an ansätze for $\rho_{\rm med}(\omega,T)$, and, in particular, a parametrization of the dominant peak. It is expected that spectral functions at nonzero temperature have a quasi-particle peak described by a Lorentzian form in the vicinity of the peak. This is confirmed by explicit calculations of the bottomonium spectral functions in the T-matrix approach Tang:2024dkz . In the T-matrix approach one calculates the in-medium scattering amplitude of heavy quark and anti-quark starting from medium-modified off-shell heavy quarks (anti-quarks) and non-perturbative interaction kernel. The non-perturbative parameters of the T-matrix approach are fixed through the vacuum quarkonium spectroscopy and comparison with lattice results on charm susceptibilities, free energy of static quarks, and Wilson line correlation functions Liu:2017qah ; Liu:2021rjf ; Tang:2023lcn ; Tang:2023tkm . Furthermore, the spectral functions corresponding to wave-function-optimized operators in the T-matrix approach can be described by a Lorentzian form around the peak, and decay exponentially far away from the peak Tang:2024dkz . As an example, in Fig. 8 we show the spectral function corresponding to the wave-function-optimized correlation function of the $\chi_{b}$(1P) state at $T=182$ MeV from the T-matrix calculations. The spectral function of a static quark anti-quark pair at high temperature calculated using hard thermal loop perturbation theory also has a similar form Burnier:2013fca . Thus, the quasi-particle peak in the spectral function can be described by the following form:
@@
with $\Theta$ being the Heaviside step function. We call this form of the spectral function the cut Lorentzian form. Such a form of the spectral function has been used in the analysis of the static quark-antiquark correlators at nonzero temperatures Bazavov:2023dci . It was shown that this form of the spectral function also leads to a nearly linear behavior of the continuum-subtracted effective masses in $\tau$ Bazavov:2023dci . In practice, a possible parametrization of $\Theta_{\delta}$ could be the following: